14 November 2014: Friday

LDAP:

  • No Issues in the past two night’s updates.
  • The Foreign Exchange Student issue seems to have resolved itself.  YAY!
  • Registrar’s office has questions about the BCA process.  Sorting that out and answering.

Backups:

  • re-created five (5) CoM/IS clients that had been deleted on Wednesday – because they are needed for their resources to be backed up.
  • Tape and CoM/IS clones completed.  Disk clones have 22TB left to go.
  • NMC does not work with Java 1.8u25 – needed to download and install the Java 1.6 from Apple.  Sigh.

FootPrints:

  • Worked through the portal help request with the developer so it stops generating emails that cause trouble for FootPrints.

Tags:

Leave a Reply

You must be logged in to post a comment.